# SUBA-Jet: a new Model for Jets in Heavy Ion Collisions

### v1

@@ V.5 Testing the role of $m_{q,g}$
Another parameter of the modeling, which varies in the different approaches, is the mass of the QGP constituents with which the jet parton scatters and which influences the various rates. For the radiative process, this mass is denoted as $m_{q}$ (with the 2 options $m_{q}=0$ or $m_{q}=\infty$). It influences the value of the radiative rate $R_{\rm rad}$ through the thermal distribution of these QGP constituents, see appendices B and C, as well as through the kinematic variables at the end of the emission of the virtual gluon. For the elastic scattering, the 2 choices we will consider here are the ‘‘energy conservation’’ and the ‘‘energy reduction’’ options, irrespective of which $m_{q}$ is taken for the inelastic collisions. By choosing the ‘‘energy conservation’’ option for the subsequent elastic scatterings of the gluon during its formation, we assume implicitly infinitely massive QGP constituents, while in the ‘‘energy reduction’’ option, the scattering is evaluated with $m_{q(g)}^{\rm therm}$ as the mass of the QGP constituents.
@@
Figure 29: Radiation spectrum of gluons from the Monte Carlo approach. The realistic radiation case with energy reduction in the elastic scattering and considering vanishing medium quark masses ($m_{q,g}=0$) (orange dashed line) is compared to the BDMPS-Z condition ($m_{q,g}=\infty$), with both energy conservation (green dotted line) and energy reduction (blue full line) prescriptions. The average number of radiated gluons per jet, $\langle N_{g}\rangle$, is listed.
Such an independence with respect to the choice of $m_{q}$ is surprising because it is not observed for the GB radiation seed, shown in Fig. 14, where a larger yield is observed for $m_{q}=0$. A finer analysis reveals that the average $\langle N_{s}\rangle$, needed to reach the critical $\varphi_{c}$, is for $\omega\gtrsim 1$ GeV larger for $m_{q}=0$ than for $m_{q}\to\infty$, as can be seen on Figs 31 and 31, leading to a larger LPM suppression. This originates from the transverse momentum given to the virtual gluon through the GB seed radiation, which is smaller for the $m_{q}=0$ case (see Fig. 15) and hence results in a smaller phase increase at early times.
@@
Figs. 34 and 34 show the momentum distribution of the gluons produced in the $m_{q/g}=\infty$cup ‘‘energy conservation’’ and in the $m_{q/g}=0$cup ‘‘energy reduction’’ scenario. They are rather different. For $m_{q/g}=0$ one sees clearly the limits due to the available phase space. The BDMPS-Z conditions ($m_{q}\to\infty$) leads even to a non-negligible backward component, which is more substantial than the one found after the GB seed (see Fig. 17). This is a consequence of the elastic collisions, which the virtual gluon suffers until it gets on-shell. These figures show that the phase space for gluons is very reduced, in longitudinal as well as in transverse direction, if we evaluate the radiation employing physical masses for the QGP partons.
@@ V.6 Testing the role of the path length $L$
Figs. 36 and 36 show the gluon energy spectrum for different path lengths in the QGP medium. Fig. 36 displays this spectrum for what we consider to be the most realistic implementation of our model (energy reduction and $m_{q}=0$). Fig. 36 shows the results for the BDMPS-Z conditions (energy conservation and $m_{q}\to\infty$). The medium path lengths varies from very small systems ($L=1$ and $2$ fm, with $\omega_{c}(L=1\text{ fm})\approx 6.25$ GeV and $\omega_{c}(L=2\text{ fm})\approx 25$ GeV) to very large systems ($L=8$ and $16$ fm, for which $\omega_{c}(L)>E$), to highlight the scaling of the radiation spectrum with respect to $L$.
We see, first of all, a large difference between the radiation spectra for the both settings. The form of the spectra for low and high $\omega$ values is rather different and at intermediate $\omega$ under the BDMPS condition roughly $10$ times more gluons are emitted than in the realistic case. This is even overcompensated by an enhanced emission of low energy gluons ($\omega\ll 1$ GeV) for the realistic case, as one can see from the average number of emitted gluons, which is in the realistic case about twice as much as for the BDMPS-Z condition. For a large path length the BDMPS-Z expectation, given by eq. (38) as $\mathrm{d}N/\mathrm{d}\omega\propto L$, should dominate the $\omega$-spectrum. This is indeed observed. Our MC results for large path lengths, $L=8$ and $16$ fm, reproduce the expected linear scaling for both settings. For a smaller path length, a reduction of the spectrum per unit length is observed, with is more pronounced for larger $\omega$ and smaller $L$, in agreement with the criteria resulting from the comparison between the formation time and the path length. We note that such a reduction is also seen at low $\omega$. This is a confirmation that this region is also impacted by coherence effects for the parameter chosen to describe the QGP ($T=400$ MeV) and by its coupling to the jet ($\alpha_{s}=0.4$). This is true for both settings although slightly more pronounced for the realistic implementation.
@@ V.7 Testing the role of gluon rescatterings and reemissions
In order to come as close as possible to the BDMPS-Z condition we performed up to now calculations with the additional assumption that collisions of formed gluons as well as elastic collisions of the jet parton are turned off. These BDMPS-Z conditions are shown in Fig. 12, where only the virtual gluon (blue-line circle) undergoes elastic scatterings with the medium and this exclusively until its formation (or discarding). No scattering of the formed gluons is allowed. The realistic case is shown in Fig. 13, where the projectile and formed gluons can have elastic collisions with the QGP and where formed gluons can emit virtual gluons themselves. Figs. 38 and 38 show the radiation spectra if we release these BDMPS-Z assumptions. Enabling emissions from gluons naturally increases the overall number of gluons per jet, but evidently only by a small amount, and especially at small $\omega$ and $k_{\perp}$ values. Only small changes in the shape of the distributions are observed, on the $\sim 10\%$ level, as shown in the bottom part of the figures.
Figure 37: Top: Gluon energy spectrum from the Monte Carlo calculation for the realistic case, taking into account elastic and inelastic scattering of formed gluons (see Fig. 13, as a dashed orange line) as compared to the BDMPS-Z conditions (see Fig. 12, shown as a solid blue line). The average number of radiated gluons per jet, $\langle N_{g}\rangle$, is listed. Bottom: Ratio of the Monte Carlo results.
Figure 38: Top: Gluon $k_{\perp}$ spectrum from the Monte Carlo calculation for the realistic case, taking into account elastic and inelastic scattering of formed gluons (shown as a dashed orange line) as compared to the BDMPS-Z conditions (shown as a solid blue line). The average number of radiated gluons per jet, $\langle N_{g}\rangle$, is listed. Bottom: Ratio of the Monte Carlo results.
@@ V.8 Global comparison of the BDMPS-Z and realistic conditions
In Figs. 40 and 40 we compare directly our results for the most realistic calculation [$m_{q}=0$, energy reduction, scattering of formed gluons and $\Delta\phi=(2P_{Q}\cdot k/E_{Q})\Delta t$] with that which comes closest to the BDMPS-Z scenario [$m_{q}\to\infty$, energy conservation, no scattering of formed gluons, $\Delta\phi=(k_{\perp}^{2}/\omega)\Delta t$]. In the top figure we display the $\omega$ and $k_{\perp}$ distribution of both approaches, in the bottom figure we present the ratio "realistic" over "BDMPS-Z like". We see that at intermediate $\omega$ both spectra have about the same functional form, however in the realistic case the yield is reduced by a factor two as compared to the BDMPS scenario. Due to the reasons discussed above the spectra differ by large factors at low and at high gluon energies. The transverse momentum spectra differ even more, in the functional form as well as in the absolute magnitude. There the conditions for the BDMPS-Z approach are too strong to allow for a realistic description of the processes which are present when a parton traverses a block of matter.
@@ VI Conclusions
In this study, we have presented a new Monte Carlo model for jet evolution and quenching in a hot and dense medium formed in ultra-relativistic heavy-ion collisions. We simulate jet evolution starting from a single hard parton of energy $E$ with maximal virtuality scale Q=E by separating two virtuality areas. In the initial high virtuality phase, Sudakov form factors are used to sample vacuum-like jet parton splittings until a jet constituent (hard parton) arrives at a low virtuality scale $Q_{0}$. The virtuality of the jet parton is modified by interactions with the medium. When the jet parton has lost its virtuality the low virtuality domain sets in. There the jet partons interact with the surrounding medium via elastic and inelastic collisions. For the latter, the radiation of virtual gluons is first sampled according to a Gunion-Bertsch approach. We have implemented the interference between the gluon emission amplitudes in the spirit of the BDMPS-Z formalism by introducing a formation phase for the virtual radiated gluons. We have employed the QCD equivalent of the LPM effect to decide whether a virtual gluon becomes finally formed.
We have presented benchmark calculations of jet energy loss in a static finite-size medium in comparison to the BDMPS-Z results. The algorithm reproduces the characteristic $\mathrm{d}N/\mathrm{d}\omega\sim\omega^{-3/2}$ behavior of the spectrum of radiated gluons when taking into account the assumptions made in the BDMPS-Z approach. For more energetic gluons it reproduces the GLV limit. For the parameters, which we have used in our studies ($\alpha_{s}=0.4$, $T=400$ MeV, $m_{g}^{\rm therm}\approx 600$ MeV), the BH regime is in fact not found at small values of $\omega\approx m_{g}^{\rm therm}$ where coherence effects are still present. In a more dilute QGP, it should however be naturally recovered in our approach.
@@
With these parameters, we have tested as well the algorithms by varying input quantities, which cannot be obtained from more fundamental approaches. They include the mass of the QGP partons, which scatter inelastically with the incoming parton and elastically with the virtual gluons, as well as the functional form of the phase accumulation of the radiation amplitude. We have also investigated different description of the elastic scattering of a gluon in the QGP medium, provided in the literature.
We have found that the spectra of gluons emitted with a large transverse momentum $k_{\perp}$ or with a large energy are not very sensitive to a variation of these quantities, with the exception of the mass of the QGP partons which influences the large $k_{\perp}$ region due to phase space constraints. At small $\omega$ and small $k_{\perp}$ the form and the absolute value of the gluon spectra depend on the choice of these input quantities. We have explained in detail the physical origin of these observations.
We have studied the path length dependence of our Monte Carlo results and found that the results are in agreement with the BDMPS-Z and GLV predictions. Finally we have relaxed all BDMPS-Z assumptions on the kinematic variables and on the included processes and have presented the gluon energy and transverse momentum spectra for what we consider to be the most realistic case. It turns out that the processes, which have been neglected in BDMPS-Z, modify the energy and transverse momentum spectra considerably. Whereas the functional form of the energy spectrum remains unchanged for intermediate energies (and therefore the $\omega^{-3/2}$ dependence of the spectra is also seen in the realistic case with, however, only half as many gluons as the BDMPS-Z approach predicts), the form of energy spectra is very different at lower and higher gluon-energies $\omega$.
@@
In a near future, we also plan to extend the benchmarking of our approach on theoretical results going beyond the soft gluon approximation and including exact $k_{\perp}$ dependence, such as Blaizot:2012fh ; Isaksen:2023nlr .
@@ Appendix A Conventions
Appendix B Sampling the Bethe-Heitler-Gunion-Bertsch seed for $m_{q}=0$
@@
In this appendix we describe the algorithm used to generate the GB pre-gluons. It is meant to be applied when the gluon energy is small with respect to the available invariant mass — which is the dominant region of radiation — while still encompassing important physical aspects — like the shrinking of the phase space — in the case of a ‘‘small’’ invariant mass. After integration over the energy- and momentum-conserving Dirac delta function the inelastic cross sections depends on five variables $x$, $\vec{k}_{\perp}$, and $\vec{l}_{\perp}$ (the momentum transferred from the light quark, transversally to the jet direction). In order to model this cross section in a transport theory one uses a Monte Carlo approach. The task is to convert a set of five uncorrelated random numbers into these variables in order to reproduce the unbiased distribution $\mathrm{d}^{5}\sigma_{\rm rad}/(\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp})$. In this appendix we describe the method presently chosen in SUBA-Jet. To draw 5 independent random numbers in the 5 dimensional $x$, $\vec{k}_{\perp}$, and $\vec{l}_{\perp}$ space and then proceed by mere rejection is very ineffective - not to say impossible - because these variables are correlated by phase space (PS) and most of the sets of random numbers would have to be discarded. To make a sampling of the cross section possible, the adopted strategy consists in finding some suitable majorant of the 5-dimensional distribution $\mathrm{d}^{5}\sigma_{\rm rad}^{m_{q}=0}/(\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp})$ that can be mapped best onto a flat 5D probabilistic-space and sufficiently close to the original distribution that the unavoidable rejection step preserves the efficiency of the sampling process by minimizing the number of trials which have to be discarded. Concretely, one first preforms recursive integrations (ascending process) on this majorant (first on $\vec{k}_{\perp}$, then on $\vec{l}_{\perp}$ and finally on $x$, resorting to additional majorations when mandatory), to obtain an invariant-mass dependent radiative cross section that is further averaged over the momentum distribution of light partons in the QGP to deduce a (majorant of the) radiative rate. The complementary (descending) process consists in sampling this rate to deduce the 3-momentum of the light parton and then, recursively, the $x$, $\vec{l}_{\perp}$ and $\vec{k}_{\perp}$ variables describing the outgoing particles. In these processes, the outgoing 3-body, [^23] The two initial partons and the radiated gluon. phase space boundary plays a specific role. Indeed, integrating the four-momentum conservation constraint results in a $1/\sqrt{\Delta(x,\vec{l}_{\perp},\vec{k}_{\perp})}$ factor in the $\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp}$ measure which leads to some superficial divergence at the PS-boundary. For finite $x$, this superficial divergence is located far beyond the ‘‘transverse’’ $(\vec{l}_{\perp},\vec{k}_{\perp})$ domain typically contributing to the radiation. Therefore, $\Delta(x,\vec{l}_{\perp},\vec{k}_{\perp})$ will be approximated to $\Delta(x,\vec{0},\vec{0})$ for practical purposes. As such an approximation may have consequences on the shrinking PS integral at small $x$$x\approx m_{g}^{2}/s$, with $s$ the invariant mass — we start by scrutinizing this case. In this appendix, we will assume that the light QGP parton is a quark of mass $m_{q}$. The generalization to scattering on gluons is pretty direct if one assumes, as we do, that it is dominated by the $t$-channel exchange. To make it easier for the reader to follow the calculation we first present a summary of the symbols, which we will use. $\sqrt{s}$ is the center of mass energy and $m_{Q},m_{q},m_{g}$ are the mass of the fast parton, of the QGP parton and of the gluon, respectively.

### v2

@@ V.5 Testing the role of $m_{q,g}$
Another parameter of the modeling, which varies in the different approaches, is the mass of the QGP constituents with which the jet parton scatters and which influences the various rates. For the radiative process, this mass is denoted as $m_{q}$ (with the 2 options $m_{q}=0$ or $m_{q}=\infty$). It influences the value of the radiative rate $R_{\rm rad}$ through the thermal distribution of these QGP constituents, see appendices B and B.4, as well as through the kinematic variables at the end of the emission of the virtual gluon. For the elastic scattering, the 2 choices we will consider here are the ‘‘energy conservation’’ and the ‘‘energy reduction’’ options, irrespective of which $m_{q}$ is taken for the inelastic collisions. By choosing the ‘‘energy conservation’’ option for the subsequent elastic scatterings of the gluon during its formation, we assume implicitly infinitely massive QGP constituents, while in the ‘‘energy reduction’’ option, the scattering is evaluated with $m_{q(g)}^{\rm therm}$ as the mass of the QGP constituents.
@@
Figure 34: Radiation spectrum of gluons from the Monte Carlo approach. The realistic radiation case with energy reduction in the elastic scattering and considering vanishing medium quark masses ($m_{q,g}=0$) (orange dashed line) is compared to that for the BDMPS-Z condition ($m_{q,g}=\infty$), with both energy conservation (green dotted line) and energy reduction (blue full line) prescriptions. The average number of radiated gluons per jet, $\langle N_{g}\rangle$, is listed.
Such an independence with respect to the choice of $m_{q}$ is surprising because it is not observed for the GB radiation seed, shown in Fig. 18, where a larger yield is observed for $m_{q}=0$. A finer analysis reveals that the average $\langle N_{s}\rangle$, needed to reach the critical $\varphi_{c}$, is for $\omega\gtrsim 1$ GeV larger for $m_{q}=0$ than for $m_{q}\to\infty$, as can be seen on Figs 36 and 36, leading to a larger LPM suppression. This originates from the transverse momentum given to the virtual gluon through the GB seed radiation, which is smaller for the $m_{q}=0$ case (see Fig. 19) and hence results in a smaller phase increase at early times.
@@
Figs. 39 and 39 show the momentum distribution of the gluons produced in the $m_{q/g}=\infty$cup ‘‘energy conservation’’ and in the $m_{q/g}=0$cup ‘‘energy reduction’’ scenario. They are rather different. For $m_{q/g}=0$ one sees clearly the limits due to the available phase space. The BDMPS-Z conditions ($m_{q}\to\infty$) leads even to a non-negligible backward component, which is more substantial than the one found after the GB seed (see Fig. 21). This is a consequence of the elastic collisions, which the virtual gluon suffers until it gets on-shell. These figures show that the phase space for gluons is very reduced, in longitudinal as well as in transverse direction, if we evaluate the radiation employing physical masses for the QGP partons.
@@
Figure 40: Primary Lund jet plane for the low virtuality shower for the kinematical BDMPS-Z conditions.
Figure 41: Same as Fig 41 but for the most realistic conditions.
In Figs. 41 and 41, we show the primary Lund plane for the low-virtuality component of our model, for the kinematical conditions of BDMPS-Z as well as for more realistic kinematical conditions. In both cases, one finds a clear structure corresponding to the growth of $k_{\perp}$ with the colinearity $1/\theta$. For increasing $\omega$, the BDMPS-Z radiation indeed satisfies the heuristic laws $\langle k_{\perp}^{2}\rangle\sim\hat{q}t_{f}\sim\sqrt{\omega\hat{q}}$ and $1/\langle\theta^{2}\rangle\sim\frac{\omega^{2}}{\langle k_{\perp}^{2}\rangle}\sim\sqrt{\omega^{3}/\hat{q}}$, which both grow with $\omega$. Passing from BDMPS-Z to more realistic conditions, one notices a reduction of the radiation at large $k_{\perp}$, already seen by comparing Figs. 39 and39, as well as an extra ultra soft radiation at large angle. In an approach, which describes the expansion of the quark gluon plasma (QGP) by hydrodynamics, this ultra soft radiation would be absorbed by the QGP.
@@ V.6 Testing the role of the path length $L$
Figs. 43 and 43 show the gluon energy spectrum for different path lengths in the QGP medium. Fig. 43 displays this spectrum for what we consider to be the most realistic implementation of our model (energy reduction and $m_{q}=0$). Fig. 43 shows the results for the BDMPS-Z conditions (energy conservation and $m_{q}\to\infty$). The medium path lengths varies from very small systems ($L=1$ and $2$ fm, with $\omega_{c}(L=1\text{ fm})\approx 6.25$ GeV and $\omega_{c}(L=2\text{ fm})\approx 25$ GeV) to very large systems ($L=8$ and $16$ fm, for which $\omega_{c}(L)>E$), to highlight the scaling of the radiation spectrum with respect to $L$.
We see, first of all, a large difference between the radiation spectra for the both settings. The form of the spectra for low and high $\omega$ values is rather different and at intermediate $\omega$ under the BDMPS condition roughly $10$ times more gluons are emitted than in the realistic case. This is even overcompensated by an enhanced emission of low energy gluons ($\omega\ll 1$ GeV) for the realistic case, as one can see from the average number of emitted gluons, which is in the realistic case about twice as large as for the BDMPS-Z condition. For a large path length the BDMPS-Z expectation, given by eq. (39) as $\mathrm{d}N/\mathrm{d}\omega\propto L$, should dominate the $\omega$-spectrum. This is indeed observed. Our MC results for large path lengths, $L=8$ and $16$ fm, reproduce the expected linear scaling for both settings. For a smaller path length, a reduction of the spectrum per unit length is observed, with is more pronounced for larger $\omega$ and smaller $L$, in agreement with the criteria resulting from the comparison between the formation time and the path length. We note that such a reduction is also seen at low $\omega$. This is a confirmation that this region is also impacted by coherence effects for the parameter chosen to describe the QGP ($T=400$ MeV) and by its coupling to the jet ($\alpha_{s}=0.4$). This is true for both settings although slightly more pronounced for the realistic implementation.
@@ V.7 Testing the role of gluon rescatterings and reemissions
In order to come as close as possible to the BDMPS-Z condition we performed up to now calculations with the additional assumption that collisions of formed gluons as well as elastic collisions of the jet parton are turned off. These BDMPS-Z conditions are shown in Fig. 16, where only the virtual gluon (blue-line circle) undergoes elastic scatterings with the medium and this exclusively until its formation (or discarding). No scattering of the formed gluons is allowed. The realistic case is shown in Fig. 17, where the projectile and formed gluons can have elastic collisions with the QGP and where formed gluons can emit virtual gluons themselves. Figs. 45 and 45 show the radiation spectra if we release these BDMPS-Z assumptions. Enabling emissions from gluons naturally increases the overall number of gluons per jet, but evidently only by a small amount, and especially at small $\omega$ and $k_{\perp}$ values. Only small changes in the shape of the distributions are observed, on the $\sim 10\%$ level, as shown in the bottom part of the figures.
Figure 44: Top: Gluon energy spectrum from the Monte Carlo calculation for the realistic case, taking into account elastic and inelastic scattering of formed gluons. The average number of radiated gluons per jet, $\langle N_{g}\rangle$, is listed. Bottom: Ratio of the Monte Carlo results.
Figure 45: Top: Gluon $k_{\perp}$ spectrum from the Monte Carlo calculation for the realistic case, taking into account elastic and inelastic scattering of formed gluons. The average number of radiated gluons per jet, $\langle N_{g}\rangle$, is listed. Bottom: Ratio of the Monte Carlo results.
@@ V.8 Global comparison of the BDMPS-Z and realistic conditions
In Figs. 47 and 47 we compare directly our results for the most realistic calculation [$m_{q}=0$, energy reduction, scattering of formed gluons and $\Delta\phi=(2P_{Q}\cdot k/E_{Q})\Delta t$] with that which comes closest to the BDMPS-Z scenario [$m_{q}\to\infty$, energy conservation, no scattering of formed gluons, $\Delta\phi=(k_{\perp}^{2}/\omega)\Delta t$]. In the top figure we display the $\omega$ and $k_{\perp}$ distribution of both approaches, in the bottom figure we present the ratio ‘‘realistic’’ over ‘‘BDMPS-Z like’’. We see that at intermediate $\omega$ both spectra have about the same functional form, however in the realistic case the yield is reduced by a factor four as compared to the BDMPS-Z scenario. Due to the reasons discussed above the spectra differ by large factors at low and at high gluon energies. The transverse momentum spectra differ even more, in the functional form as well as in the absolute magnitude. There the conditions for the BDMPS-Z approach are too strong to allow for a realistic description of the processes which are present when a parton traverses a block of matter.
@@ VI Conclusions
In this study, we have presented a new Monte Carlo model for jet evolution and quenching in a hot and dense medium formed in ultra-relativistic heavy-ion collisions. We simulate jet evolution starting from a single hard parton of energy $E$ with a maximal virtuality scale $Q=E$ by separating two virtuality areas. In the initial high virtuality phase, Sudakov form factors are used to sample vacuum-like jet parton splittings until a jet constituent (hard parton) arrives at a low virtuality scale $Q_{0}$. The virtuality of the jet parton is modified by interactions with the medium. When the jet parton has lost its virtuality the low virtuality domain sets in. There the jet partons interact with the surrounding medium via elastic and inelastic collisions. For the latter, the radiation of virtual gluons is first sampled according to a Gunion-Bertsch approach. We have implemented the interference between the gluon emission amplitudes, specific to the BDMPS-Z formalism, in the spirit of Zapp:2011ya , by introducing a formation phase for the virtual radiated gluons. We took into account the exact phase space limitation what allowed us to explore the whole ($\omega,k_{T}$) space. This has several consequences: a) we can go beyond the eikonal approximation and explore the full kinematic range encountered in the study of an expanding plasma, b) we can study the recoil given to the QGP parton what allows a feed back of the kinematical variables of the jet-QGP interaction to the QGP. c) we can explore the gluon energy distribution beyond the range of validity of the BDMPS-Z approach and recover for larger values of $\omega$ the spectral form predicted by GLV and d) we can study the influence of a finite mass of the QGP parton on the gluon spectra.
We have presented benchmark calculations of jet energy loss in a static finite-size medium in comparison to the BDMPS-Z results. The algorithm reproduces the characteristic $\mathrm{d}N/\mathrm{d}\omega\sim\omega^{-3/2}$ behavior of the spectrum of radiated gluons when taking into account the assumptions made in the BDMPS-Z approach. For more energetic gluons we reproduce the GLV limit. For the parameters, which we have used in our studies ($\alpha_{s}=0.4$, $T=400$ MeV, $m_{g}^{\rm therm}\approx 600$ MeV), the BH regime is in fact not found at small values of $\omega\approx m_{g}^{\rm therm}$ where coherence effects are still present. In a more dilute QGP, it should however be naturally recovered in our approach.
@@
With these parameters, we have tested as well the algorithms by varying input quantities, which cannot be obtained from more fundamental approaches. They include the mass of the QGP partons, which scatter inelastically with the incoming parton and elastically with the virtual gluons, as well as the functional form of the phase accumulation of the radiation amplitude. We have also investigated different descriptions of the elastic scattering of a gluon in the QGP medium, provided in the literature.
We have found that the spectra of gluons emitted with a large transverse momentum $k_{\perp}$ or with a large energy are not very sensitive to a variation of these quantities, with the exception of the mass of the QGP partons, which influences the large $k_{\perp}$ region due to phase space constraints. At small $\omega$ and small $k_{\perp}$ the form and the absolute value of the gluon spectra depend on the choice of these input quantities. We have explained in detail the physical origin of these observations.
We have studied the path length dependence of our Monte Carlo results and found that the results are in agreement with the BDMPS-Z and GLV predictions. Finally we have relaxed all BDMPS-Z assumptions on the kinematic variables and on the included processes and have presented the gluon energy and transverse momentum spectra for what we consider to be the most realistic case. It turns out that the processes, which have been neglected in BDMPS-Z, modify the energy and transverse momentum spectra considerably. Whereas the functional form of the energy spectrum remains unchanged for intermediate energies (and therefore the $\omega^{-3/2}$ dependence of the spectra is also seen in the realistic case with, however, only a quarter of the gluons as compared to the BDMPS-Z prediction), the form of energy spectra is very different at lower and higher gluon-energies $\omega$.
@@
In a near future, we also plan to extend the benchmarking of our approach on theoretical results going beyond the small x- approximation and including the exact $k_{\perp}$ dependence, such as Blaizot:2012fh ; Isaksen:2023nlr .
@@ Appendix A Conventions
Appendix B Sampling of the seed for the Bethe-Heitler-Gunion-Bertsch gluon emission
In this appendix we describe the algorithm used to generate the GB pre-gluons. It is valid in the dominant region of radiation where the gluon energy is small with respect to the available invariant mass. In the opposite case of a small invariant mass, the exact measure is slightly simplified for the sake of the effectiveness of the MC sampling, which however still encompasses important physical aspects — like the shrinking of the phase space.
@@ Appendix B Sampling of the seed for the Bethe-Heitler-Gunion-Bertsch gluon emission
After integration over the energy- and momentum-conserving Dirac delta function, the inelastic cross sections depends on five variables $x$, $\vec{k}_{\perp}$, and $\vec{l}_{\perp}$ (the momentum transferred from the light quark, transversally to the jet direction). In order to model this cross section in a transport theory one uses a Monte Carlo approach. The task is to convert a set of five uncorrelated random numbers into these variables in order to reproduce the unbiased distribution $\mathrm{d}^{5}\sigma_{\rm rad}/(\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp})$. In this appendix we describe the method presently chosen in SUBA-Jet. To draw 5 independent random numbers in the 5 dimensional $x$, $\vec{k}_{\perp}$, and $\vec{l}_{\perp}$ space and then proceed by mere rejection is very ineffective - not to say impossible - because these variables are correlated by phase space (PS) and most of the sets of random numbers would have to be discarded. To make a sampling of the cross section possible, the adopted strategy consists in finding some suitable majorant of the 5-dimensional distribution $\mathrm{d}^{5}\sigma_{\rm rad}/(\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp})$ that can be mapped best onto a flat 5D probabilistic-space and sufficiently close to the original distribution that the unavoidable rejection step preserves the efficiency of the sampling process by minimizing the number of trials, which have to be discarded. Concretely, one first preforms recursive integrations (ascending process) on this majorant (first on $\vec{k}_{\perp}$, then on $\vec{l}_{\perp}$ and finally on $x$, resorting to additional majorations when mandatory), to obtain an invariant-mass dependent radiative cross section that is further averaged over the momentum distribution of light partons in the QGP to deduce a (majorant of the) radiative rate. The complementary (descending) process consists in sampling this rate to deduce the 3-momentum of the light parton and then, recursively, the $x$, $\vec{l}_{\perp}$ and $\vec{k}_{\perp}$ variables describing the outgoing particles. In these processes, the outgoing 3-body, [^24] The two initial partons and the radiated gluon. phase space boundary plays a specific role. Indeed, integrating the four-momentum conservation constraint results in a $1/\sqrt{\Delta(x,\vec{l}_{\perp},\vec{k}_{\perp})}$ factor in the $\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp}$ measure which leads to some superficial divergence at the PS-boundary. For finite $x$, this superficial divergence is located far beyond the ‘‘transverse’’ $(\vec{l}_{\perp},\vec{k}_{\perp})$ domain typically contributing to the radiation. Therefore, $\Delta(x,\vec{l}_{\perp},\vec{k}_{\perp})$ will be approximated to $\Delta(x,\vec{0},\vec{0})$ for practical purposes. As such an approximation may have consequences on the shrinking PS integral at small $x$$x\approx m_{g}^{2}/s$, with $s$ the invariant mass — we start by scrutinizing this case. We present the majorant to the radiation cross section for two limits:
a) for the assumption that the mass of the medium quark is finite, restricting ourselves to $m_{q}=0$ for the concrete expressions: $\mathrm{d}^{5}\sigma_{\rm rad}^{m_{q}=0}/(\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp})$ and
b) for the assumption that the mass $m_{q}\to\infty$: $\mathrm{d}^{5}\sigma_{\rm rad}^{m_{q}\to\infty}/(\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp})$
In this appendix, we will assume that the light QGP parton is a quark of mass $m_{q}$. The generalization to scattering on gluons is pretty direct if one assumes, as we do, that it is dominated by the $t$-channel exchange.
B.1 Definition of the symbols
To make it easier for the reader to follow the calculation we first present a summary of the symbols, which we will use. $\sqrt{s}$ is the center of mass energy and $m_{Q}$, $m_{q}$, $m_{g}$ are the mass of the fast parton, of the QGP parton and of the gluon, respectively.